OpenStudy (anonymous):

solving quadratics using the quadratic formula Solve. (x – 5)^2 = 2

hero (hero):

You don't need quadratic formula to solve this

OpenStudy (anonymous):

but it wants us to use the quadratic formula to solve it @Hero

hero (hero):

Just simply take the square root of both sides, then add 5 to both sides. That's all you have to do in order to solve for x. Using the quadratic formula is pointless in this case.

OpenStudy (anonymous):

ok thanks @Hero

OpenStudy (anonymous):

Solve. x2 = 6x – 6 @hero do you think you can help with this one as well i dont want the answer i just want help

hero (hero):

(x – 5)^2 = 2 Square root both sides: x - 5 = ±sqrt(2) x = ±sqrt(2) + 5

OpenStudy (anonymous):

okay i got it right then thanks now can you help with the second not the answer just help @Hero

hero (hero):

x2 = 6x – 6 x^2 - 6x - 6 = 0 In this case, use the quadratic formula: a = 1, b = -6, c = -6

hero (hero):

Quadratic Formula: \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\]

hero (hero):

Plug and play

OpenStudy (anonymous):

is the answer \[X=3\pm \sqrt{3}\] @Hero

hero (hero):

\[x = \frac{-(-6) \pm \sqrt{(-6)^2 - 4(1)(-6)}}{2(1)}\]

hero (hero):

\[x = \frac{6 \pm \sqrt{36 +24}}{2}\] \[x = \frac{6 \pm \sqrt{60}}{2}\] \[x = 3 \pm \sqrt{60}\] \[x = 3 \pm \sqrt{4 \dot\ 15}\] \[x = 3 \pm \sqrt{4}\sqrt{15}\] \[x = 3 \pm 2\sqrt{15}\]

OpenStudy (anonymous):

what aboutthis one What are the x– and y–intercepts of y = 2x2 – 10x + 8?

OpenStudy (anonymous):

@hero sorry for bothering you so much

hero (hero):

I'm pretty sure you can figure this one out on your own

OpenStudy (anonymous):

yeah i just answered it lol thanks @Hero

hero (hero):

To find x intercept, let y = 0 and solve for x. To find y intercept, let x = 0 and solve for y

OpenStudy (anonymous):

yeah thanks :) @Hero